OpenStudy (sleepyjess):

Factor the algebraic expression in terms of a single trigonometric function. \(\large cos~x~ -~ sin^2x - 1\)

OpenStudy (sleepyjess):

@phi

OpenStudy (phi):

i think they want you to use sin^2 + cos^2 = 1 to get rid of the sin^2 (replace it with 1 - cos^2 )

OpenStudy (sleepyjess):

So, \(\cos x + 1 - \cos^2\)?

OpenStudy (phi):

close. check your algebra

OpenStudy (phi):

cos x - (1 - cos^2 x ) - 1

OpenStudy (phi):

distribute the -1. after you simplify, you have a "quadratic" that you can factor.

OpenStudy (sleepyjess):

cosx-1+cos^2 x -1?

OpenStudy (phi):

ok, but you can combine -1 - 1 and it's always good to put it in "standard order" so make cos^2 x the first term

OpenStudy (sleepyjess):

Ok, \(\cos^2x+\cos x-2\)

OpenStudy (phi):

yes. now let's rename cos(x) as y. this makes the expression \[ y^2 + y -2 \] can you factor that ? (yes you can!)

OpenStudy (phi):

(y + ) (y - ) look at the 2 and list its factors: 1,2 (only choice) they have different signs, and the larger one will be +

OpenStudy (phi):

you should get (y+2)(y-1) or, putting cos x back in (cos x +2)( cos x -1)

OpenStudy (sleepyjess):

Thank you!